OpenStudy (anonymous):

Calculus U-substition HELP PLEASE :) Integral 1/3 sec^8x dx Thanks!

zepdrix (zepdrix):

Is the secant in the denominator?

OpenStudy (anonymous):

nope numerator

zepdrix (zepdrix):

sec^8x = sec^2x*sec^6x = sec^2x(sec^2x)^3 = sec^2x(tan^2x+1)^3 Understand what I did there? :U From here we can apply a nice easy substitution. I won't spill the beans if you want to find it from this step. Lemme know if you're still stuck.